# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 1 Introduction
The recognition and interpretation of hand gestures (ADADOGLOU) play a central role in advancing Human–Computer Interaction (HCI) and assistive communication systems. In recent years, deep learning–based visual hand gesture recognition (VHGR) has gained significant momentum as an active subfield of computer vision, building upon progress in related areas such as action (survey-HAR-introduction-shafizadegan2024multimodal) and facial expression recognition (survey-FER-JESWA-2024). This has led to increased focus on Hand Gesture Recognition (HGR) across various applications. For example, HGR is widely used in Human-Robot Interaction (HRI) in industrial settings (e.g. robotic arm control (Papadopoulos2021Access; ROBOTIC_ARM)), in healthcare (e.g. robot-assisted surgery (TELEOPERATED_SURGICAL_ROBOT)), and in the autonomous navigation of mobile robots such as Unmanned Aerial (MD-UHGRD_SA-YOLO) and Ground (ULTRA-RANGE-UGV-Bamani; Linardakis2024Access) Vehicles (UAVs and UGVs). It also plays a key role in HCI, especially in AR/VR environments (OO-dMVMT; EgoGesture_DATASET; BOAT-MI_VR). In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic Sign Language Recognition (SLR) systems, aiming to improve communication and inclusion for these communities (survey-deaf-mute-AIR-2024; NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ).
@@
(ADADOGLOU) 2016-2020 - Vision-based DL CSLR Provides an in-depth comparison of existing approaches, loss functions and pretraining schemes for CSLR. This survey expands the scope including non-CSLR methods and examines more recent methods.
@@
(survey-5-9622242) 2014-2020 98 Vision-based DL & ML General Investigates the progress on hand gesture representation and Data Acquisition (DAQ) techniques as well as their performance and future research directions. It examines relatively antiquated methods. The current work aims to emphasize on recent DL approaches and related challenges and future perspectives.
@@
(survey-10-al2021deep) 2014-2021 138 Sensor- & Vision-based DL SLR Categorizes Continuous and Isolated SLR methods based on their architectures and discusses pose estimation for ISLR in detail. This survey covers more topics including HCI & HRI, and highlights task-specific architectural patterns and challenges for CSLR.
@@
(survey-11-adeyanju2021machine) 2001-2021 649 Vision-based DL & ML SLR Discusses relationships between research teams and DAQ techniques. It includes relatively old methods. This survey emphasizes on recently proposed ones in order to facilitate the evolution of the field.
@@
(survey-6-ojeda2022survey) 2000-2020 571 Sensor- & Vision-based DL & ML General Provides many statistical results and bibliometric charts. It does not identify contemporary challenges or future research directions; this work seeks to detect relevant material.
@@
(survey-3-el2022comprehensive) 1998-2020 - Sensor- & Vision-based DL & ML SLR Discusses in detail existing methods for SLR and provides multiple tables for quantitative comparison. This survey analyzes newer approaches for SLR and expands taxonomy and comparison to more application domains, beyond sign language.
@@
(survey-16-joksimoski2022technological) 2010-2021 80 Vision-based DL & ML SLR A Natural Language Processing (NLP) toolkit was utilized to filter publications. Insightful statistics are also provided. Only a few methods are discussed and the current trends are not highlighted. On the contrary, this survey aims to identify trends and challenges.
@@
(survey-17-sarhan2023unraveling) 2015-2023 - Vision-based DL ISLR Provides statistical charts and summary tables based on a four-fold taxonomy. This review introduces additional taxonomy criteria and covers more tasks.
@@
(survey-4-shin2024methodological) 2014-2024 - Sensor- & Vision-based DL & ML General Considers a variety of modalities and examines methods for each one. The current work aims to discuss further multi-modality fusion strategies and identify more challenges.
@@
(survey-8-hashi2024systematic) 2018-2023 256 Sensor- & Vision-based ML General Covers various aspects of HGR, including DAQ techniques, hand segmentation and feature extraction. Focuses exclusively on traditional ML approaches, which are currently surpassed by DL ones. This survey, searching for recent advancements, focuses on DL approaches.
@@
(survey-9-al2024advancements) 2015-2024 58 Sensor- & Vision-based DL & ML SLR Examines various SLR methods and highlights the superiority of DL methods over older ones. Its does not identify challenges and future perspectives; this survey aims to cover this, going deeper into details.
@@
(survey-13-tao2024sign) 2017-2024 - Sensor- & Vision-based DL & ML SLR Groups methods into RNN-based, CNN-based and transformers whilst discusses well-known benchmarks. This survey expands the scope to non-SLU papers and seeks to provide methods comparison and future prospects.
@@
(survey-1-rahman2024comparative) 2008-2024 - Sensor- & Vision-based DL & ML General Covers a wide spectrum of gesture-based interaction and . The considered methods are relatively old. This survey aims to identify more recent advancements.
@@
(linardakis2025_hgr_survey) 2018-2025 125 Vision-based DL & ML General Discuss jointly hand gesture capturing and recognition techniques. Many challenges are highlighted, including data bias and privacy issues. It covers a limited number of datasets for HGR and relatively old methods; this survey complements it by analyzing more datasets and recent advancements.
@@
(survey-emporio2025continuous) 2002-2023 - Sensor- & Vision-based DL & ML CDHGR Focuses on online gesture recognition, discussing DAQ technologies and evaluation metrics, among others. This survey covers more topics and provides comparison of state-of-the-art methods.
@@
(survey-khan2025deep) 2010-2024 - Sensor- & Vision-based DL & ML CSLR Discusses in detail methods and datasets for CSLR. It does not conduct methods comparison and the challenges are general; this survey highlights optimization difficulties in the context of CSLR.
@@
The great importance and wide scope of HGR has led researchers to conduct surveys focusing on different aspects of the field, as shown in Table 1. Compared to existing review papers, this survey provides an in-depth examination of the recent Deep Learning (DL) methods (published since 2021), covering the current state-of-the-art (SoTA). In contrast, other existing reviews either focus on older approaches (published since 1998) (survey-6-ojeda2022survey; survey-3-el2022comprehensive; survey-8-hashi2024systematic) or mainly cover traditional machine learning techniques (survey-1-rahman2024comparative; survey-13-tao2024sign; linardakis2025_hgr_survey). While such works are valuable for understanding the fundamental methods and providing statistical insights, they are less suited to reflecting the recent evolution of the field. The present survey stands out by its broader and more up-to-date scope, including 238 papers published since 2021—-an average of nearly 47 publications per year. By comparison, (survey-11-adeyanju2021machine) and (survey-6-ojeda2022survey) analyze 649 and 571 papers respectively, but span roughly 20 years of research, corresponding to only about 30 publications per year. Finally, the current survey focuses exclusively on vision-based HGR. Including sensor-based approaches would significantly expand the scope and potentially confuse readers, while such methods generally offer limited expressiveness and reduced applicability in real-world scenarios. Nevertheless, the review maintains a broad and inclusive perspective, acknowledging that approaches from different domains may share common principles and can inform each other.
@@
More specifically, the survey focuses on vision-based methods for three key VHGR tasks, which are discussed in detail in the following sections: static HGR (SHGR) (FGDSNet; HAGRIDv2_ARXIV), isolated dynamic HGR (IDHGR) (SAM_SLR; SignBERT_ICCV; NLA_SLR; CTFB_Hampiholi; RAAR3DNet), and continuous (dynamic) HGR (CDHGR or CHGR) (VAC-CSLR; NEURIPS2022_6cd3ac24_HEATMAPS; CorrNet). SHGR involves recognizing gestures from single frames or spatial data only. IDHGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns. CDHGR goes a step further by detecting multiple gestures within long, unsegmented video sequences. Each task presents its own unique challenges. For example, CHGR often relies on weak supervision, making optimization more difficult. At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
@@
VHGR has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines (NVIDIA_GESTURE_R3DCNN_Molchanov). Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification. Pre-processing often involved segmenting the hand region using skin color or other visual cues. Feature extraction relied on descriptors like Histogram of Oriented Gradients (HOG), while classification was usually performed using models such as Support Vector Machines (SVMs) or Hidden Markov Models (HMMs), depending on the task. With the rise of deep learning, these hand-crafted features were gradually replaced by learned representations, which offer better generalization. Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks. These models first capture short-term motion patterns and then aggregate them to understand long-term context (VAC-CSLR; TemporalLiftPooling; SMKD; SEN-CSLR; CorrNet). Due to their superior performance, this study focuses primarily on deep learning-based methods.
@@ 3.1 Application Domain
    Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU (UNISIGNli2025; BOBSL-BBC)) or across multiple sign languages (multilingual SLU (SIGN-muller-etal-2023-findings)).
@@
    Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers. Numerous studies focus on HCI (HAGRIDv2_ARXIV; ZJUGesture), with particular emphasis on AR/VR applications (EgoGesture_DATASET; OO-dMVMT; BOAT-MI_VR).
@@
    Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs (ULTRA-RANGE-UGV-Bamani; MOBILE_ROBOT_DRCAM) and UAVs (LONG_RANGE_HGR_SSD; MD-UHGRD_SA-YOLO), industrial robotic arms (ROBOTIC_ARM), assistive robots (DIALOG_POINTING), and surgical robots (TELEOPERATED_SURGICAL_ROBOT).
@@
    Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts. VHGR enables recognition of traffic gestures (e.g., from police officers) (FA-STGCN; MG-GCT; CTCX_WANG2022123) and supports the development of more intuitive in-vehicle user interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; EgoFormer_METHOD_EgoDriving_DATASET; ConvMixFormer; Briareo_dataset; MVTN_ARXIV).
@@ 3.2 Task
Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics. Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time. Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction (NVIDIA_GESTURE_R3DCNN_Molchanov), which reflect the gesture’s evolution from initiation to completion. The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
@@ 3.3 Input
In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR. The choice of input modality directly influences the representation architecture and the deep learning models employed. While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps (including RGB-like representations, such as HHA (SAM_SLR)), infrared (IR) images, skeleton joints (and derived features like bones and motion (TMS-NET)), pose heatmaps (NEURIPS2022_6cd3ac24_HEATMAPS; NLA_SLR), optical flow and motion history (MOTION_HISTORY_Mercanoglu).
In terms of the number of input modalities, methods are classified as uni-modal, when they use a single type of input data, or multi-modal, when they integrate two or more modalities. In the case of multiple modalities in the input, a multi-modality fusion strategy is chosen (MMTM; CTFB_Hampiholi). This could be:
@@
    Early (data-level) fusion: Combines modalities at the input level before processing (DHGR_3D_POSE_HRI; 3STREAM_DISTILLATION).
@@
    Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models (MOTION_HISTORY_Mercanoglu; JOINTS_MOTION_VECTORS_Improved_ST-GCN; GestFormer).
@@
    Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  (tang2021graph).
@@
    Multi-level fusion: Enables interaction across modalities at multiple stages of processing (NLA_SLR; CTFB_Hampiholi).
@@
(a) From left to right: RGB, Optical Flow, Depth and IR (NVIDIA_GESTURE_R3DCNN_Molchanov)
(b) Skeleton joints (UNISIGNli2025)
(c) Pose heatmap (NEURIPS2022_6cd3ac24_HEATMAPS)
(d) Motion history (MOTION_HISTORY_Mercanoglu)
@@ 3.4 Type of Information Conveyed
Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures. Specifically, it examines the kind of information gestures are meant to transmit (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET), resulting in the following categories (illustrated in Fig. 4):
@@
    Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; HAGRIDv2_ARXIV; ULTRA-RANGE-UGV-Bamani).
@@
    Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages (BOBSL-BBC; Auslan-Daily).
@@
    Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets (DeePoint).
@@
    Manipulative gestures aim to interact with or mimic the manipulation of physical objects (FPHA).
(a) Semaphoric (WACV_kapitanov2024hagrid)
(b) Language (BOBSL-BBC)
(c) Pointing (DeePoint)
(d) Manipulative (FPHA)

### v3

@@ 1 Introduction
The recognition and interpretation of hand gestures (2) play a central role in advancing Human–Computer Interaction (HCI) and assistive communication systems. In recent years, deep learning–based visual hand gesture recognition (VHGR) has gained significant momentum as an active subfield of computer vision, building upon progress in related areas such as action (175) and facial expression recognition (169). This has led to increased focus on Hand Gesture Recognition (HGR) across various applications. For example, HGR is widely used in Human-Robot Interaction (HRI) in industrial settings (e.g. robotic arm control (151; 107)), in healthcare (e.g. robot-assisted surgery (156)), and in the autonomous navigation of mobile robots such as Unmanned Aerial (213) and Ground (18; 116) Vehicles (UAVs and UGVs). It also plays a key role in HCI, especially in AR/VR environments (34; 220; 5). In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic Sign Language Recognition (SLR) systems, aiming to improve communication and inclusion for these communities (216; 27).
@@
(2) 2016-2020 - Vision-based DL CSLR Provides an in-depth comparison of existing approaches, loss functions and pretraining schemes for CSLR. This survey expands the scope including non-CSLR methods and examines more recent methods.
@@
(137) 2014-2020 98 Vision-based DL & ML General Investigates the progress on hand gesture representation and Data Acquisition (DAQ) techniques as well as their performance and future research directions. It examines relatively antiquated methods. The current work aims to emphasize on recent DL approaches and related challenges and future perspectives.
@@
(7) 2014-2021 138 Sensor- & Vision-based DL SLR Categorizes Continuous and Isolated SLR methods based on their architectures and discusses pose estimation for ISLR in detail. This survey covers more topics including HCI & HRI, and highlights task-specific architectural patterns and challenges for CSLR.
@@
(3) 2001-2021 649 Vision-based DL & ML SLR Discusses relationships between research teams and DAQ techniques. It includes relatively old methods. This survey emphasizes on recently proposed ones in order to facilitate the evolution of the field.
@@
(146) 2000-2020 571 Sensor- & Vision-based DL & ML General Provides many statistical results and bibliometric charts. It does not identify contemporary challenges or future research directions; this work seeks to detect relevant material.
@@
(46) 1998-2020 - Sensor- & Vision-based DL & ML SLR Discusses in detail existing methods for SLR and provides multiple tables for quantitative comparison. This survey analyzes newer approaches for SLR and expands taxonomy and comparison to more application domains, beyond sign language.
@@
(93) 2010-2021 80 Vision-based DL & ML SLR A Natural Language Processing (NLP) toolkit was utilized to filter publications. Insightful statistics are also provided. Only a few methods are discussed and the current trends are not highlighted. On the contrary, this survey aims to identify trends and challenges.
@@
(174) 2015-2023 - Vision-based DL ISLR Provides statistical charts and summary tables based on a four-fold taxonomy. This review introduces additional taxonomy criteria and covers more tasks.
@@
(181) 2014-2024 - Sensor- & Vision-based DL & ML General Considers a variety of modalities and examines methods for each one. The current work aims to discuss further multi-modality fusion strategies and identify more challenges.
@@
(71) 2018-2023 256 Sensor- & Vision-based ML General Covers various aspects of HGR, including DAQ techniques, hand segmentation and feature extraction. Focuses exclusively on traditional ML approaches, which are currently surpassed by DL ones. This survey, searching for recent advancements, focuses on DL approaches.
@@
(6) 2015-2024 58 Sensor- & Vision-based DL & ML SLR Examines various SLR methods and highlights the superiority of DL methods over older ones. Its does not identify challenges and future perspectives; this survey aims to cover this, going deeper into details.
@@
(192) 2017-2024 - Sensor- & Vision-based DL & ML SLR Groups methods into RNN-based, CNN-based and transformers whilst discusses well-known benchmarks. This survey expands the scope to non-SLU papers and seeks to provide methods comparison and future prospects.
@@
(157) 2008-2024 - Sensor- & Vision-based DL & ML General Covers a wide spectrum of gesture-based interaction and . The considered methods are relatively old. This survey aims to identify more recent advancements.
@@
(117) 2018-2025 125 Vision-based DL & ML General Discuss jointly hand gesture capturing and recognition techniques. Many challenges are highlighted, including data bias and privacy issues. It covers a limited number of datasets for HGR and relatively old methods; this survey complements it by analyzing more datasets and recent advancements.
@@
(47) 2002-2023 - Sensor- & Vision-based DL & ML CDHGR Focuses on online gesture recognition, discussing DAQ technologies and evaluation metrics, among others. This survey covers more topics and provides comparison of state-of-the-art methods.
@@
(99) 2010-2024 - Sensor- & Vision-based DL & ML CSLR Discusses in detail methods and datasets for CSLR. It does not conduct methods comparison and the challenges are general; this survey highlights optimization difficulties in the context of CSLR.
@@
The great importance and wide scope of HGR has led researchers to conduct surveys focusing on different aspects of the field, as shown in Table 1. Compared to existing review papers, this survey provides an in-depth examination of the recent Deep Learning (DL) methods (published since 2021), covering the current state-of-the-art (SoTA). In contrast, other existing reviews either focus on older approaches (published since 1998) (146; 46; 71) or mainly cover traditional machine learning techniques (157; 192; 117). While such works are valuable for understanding the fundamental methods and providing statistical insights, they are less suited to reflecting the recent evolution of the field. The present survey stands out by its broader and more up-to-date scope, including 238 papers published since 2021—-an average of nearly 47 publications per year. By comparison, (3) and (146) analyze 649 and 571 papers respectively, but span roughly 20 years of research, corresponding to only about 30 publications per year. Finally, the current survey focuses exclusively on vision-based HGR. Including sensor-based approaches would significantly expand the scope and potentially confuse readers, while such methods generally offer limited expressiveness and reduced applicability in real-world scenarios. Nevertheless, the review maintains a broad and inclusive perspective, acknowledging that approaches from different domains may share common principles and can inform each other.
@@
More specifically, the survey focuses on vision-based methods for three key VHGR tasks, which are discussed in detail in the following sections: static HGR (SHGR) (229; 145), isolated dynamic HGR (IDHGR) (91; 78; 235; 66; 227), and continuous (dynamic) HGR (CDHGR or CHGR) (136; 29; 84). SHGR involves recognizing gestures from single frames or spatial data only. IDHGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns. CDHGR goes a step further by detecting multiple gestures within long, unsegmented video sequences. Each task presents its own unique challenges. For example, CHGR often relies on weak supervision, making optimization more difficult. At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
@@
VHGR has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines (138). Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification. Pre-processing often involved segmenting the hand region using skin color or other visual cues. Feature extraction relied on descriptors like Histogram of Oriented Gradients (HOG), while classification was usually performed using models such as Support Vector Machines (SVMs) or Hidden Markov Models (HMMs), depending on the task. With the rise of deep learning, these hand-crafted features were gradually replaced by learned representations, which offer better generalization. Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks. These models first capture short-term motion patterns and then aggregate them to understand long-term context (136; 83; 69; 85; 84). Due to their superior performance, this study focuses primarily on deep learning-based methods.
@@ 3.1 Application Domain
    Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU (113; 12)) or across multiple sign languages (multilingual SLU (140)).
@@
    Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers. Numerous studies focus on HCI (145; 212), with particular emphasis on AR/VR applications (220; 34; 5).
@@
    Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs (18; 172) and UAVs (232; 213), industrial robotic arms (107), assistive robots (31), and surgical robots (156).
@@
    Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts. VHGR enables recognition of traffic gestures (e.g., from police officers) (50; 65; 204) and supports the development of more intuitive in-vehicle user interfaces (138; 155; 58; 128; 57).
@@ 3.2 Task
Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics. Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time. Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction (138), which reflect the gesture’s evolution from initiation to completion. The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
@@ 3.3 Input
In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR. The choice of input modality directly influences the representation architecture and the deep learning models employed. While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps (including RGB-like representations, such as HHA (91)), infrared (IR) images, skeleton joints (and derived features like bones and motion (40)), pose heatmaps (29; 235), optical flow and motion history (133).
In terms of the number of input modalities, methods are classified as uni-modal, when they use a single type of input data, or multi-modal, when they integrate two or more modalities. In the case of multiple modalities in the input, a multi-modality fusion strategy is chosen (195; 66). This could be:
@@
    Early (data-level) fusion: Combines modalities at the input level before processing (54; 198).
@@
    Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models (133; 186; 56).
@@
    Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  (190).
@@
    Multi-level fusion: Enables interaction across modalities at multiple stages of processing (235; 66).
@@
(a) From left to right: RGB, Optical Flow, Depth and IR (138)
(b) Skeleton joints (113)
(c) Pose heatmap (29)
(d) Motion history (133)
@@ 3.4 Type of Information Conveyed
Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures. Specifically, it examines the kind of information gestures are meant to transmit (17), resulting in the following categories (illustrated in Fig. 4):
@@
    Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces (138; 145; 18).
@@
    Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages (12; 177).
@@
    Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets (141).
@@
    Manipulative gestures aim to interact with or mimic the manipulation of physical objects (55).
(a) Semaphoric (95)
(b) Language (12)
(c) Pointing (141)
(d) Manipulative (55)